OpenStudy (anonymous):

Find the standard matrix for the orthogonal projection of R^3 onto the span{a_1,a_2} and confirm that the matrix is idempotent? a_1=(3, -4, 1) a_2=(2,0,3)

OpenStudy (zarkon):

\[\left[\begin{matrix}\frac{113}{257}&\frac{-84}{257} &\frac{96}{257} \\ \frac{-84}{257} & \frac{208}{257} &\frac{56}{257} \\\frac{96}{257} & \frac{56}{257} &\frac{193}{257} \end{matrix}\right]\]

OpenStudy (anonymous):

Yeah that is the answer! how'd you get it though??

OpenStudy (zarkon):

I found an orthogonal basis for the span then I found the projection matrix for each basis vector of my span then I added the two matrices

OpenStudy (anonymous):

I can't get the answer for some reason.. did you use 1/a^t*a (a*a^T) for each vector?

OpenStudy (anonymous):

then add them together sorry I meant a^T

OpenStudy (zarkon):

did you find an orthogonal basis first

OpenStudy (zarkon):

you should use \[\frac{v^Tv}{v\cdot v}\]

OpenStudy (zarkon):

http://alturl.com/c5xiq

OpenStudy (zarkon):

http://alturl.com/6ez98

OpenStudy (anonymous):

Oh you used the gram schmidt process

OpenStudy (anonymous):

okay thanks